OpenStudy (anonymous):

The hypotenuse of a right triangle measures 13 centimeters. find the lengths of the legs if their sum is 17 centimeters

OpenStudy (solomonzelman):

\[a^2+b^2=c^2\] c is the hypotenuse. a and b are the legs. Say that leg a is x and leg b is 17-x. Can you find the legs now? So you have \[x^2 + (17-x)^2=13^2\] More help?

OpenStudy (solomonzelman):

\[x^2+289-34x+x^2= 169\] \[2x^2-34x+289=169\] \[2x^2-34x=169-289\] \[2x^2-34x=-120\] or you can have 120 on the other side, \[2x^2-34x+120=0\] factor the equation out of 2, in other words divide entire equation (everything in it) by 2. \[x^2-17x+60=0\] \[Can+You+Solve+It+Now+?\]

OpenStudy (anonymous):

yes, thank you it took a few min.